OpenStudy (anonymous):

combining like terms N- 4- 9 I got -13n

OpenStudy (anonymous):

n-13

OpenStudy (anonymous):

n-13=0 so, therefore, n=13

OpenStudy (anonymous):

No, it's not n=13

OpenStudy (anonymous):

show me how nis not = to 13?

OpenStudy (anonymous):

becuase they didnt say that it equals to zero.. or a number or whatever... so you can't solve it!

OpenStudy (turingtest):

it doesn't say anything equals zero, so all we can do is put N-13 is the simplified form of the above

OpenStudy (turingtest):

the answer is N-13 in simplified form

OpenStudy (anonymous):

^ exactly :)

OpenStudy (anonymous):

true, if they are not asking for it to be solved.

OpenStudy (anonymous):

ok my first answer was -13n so I need to keep the n in front

OpenStudy (anonymous):

it says simplyfy each expression

OpenStudy (anonymous):

Question Turing I was always taught that the letter goes on the right. What makes this different?

OpenStudy (turingtest):

what you need to to is only mess with the numbers that have like terms: examples: 5a+6a=11a (because they both have an "a" we can combine them) 5x-12+6=5x+6 (we have to put together the parts that have no x) 2x+3y+4x+5y=(2x+4x)+(3y+5y)=6x+8y This last example really shows what combining like terms is all about: putting the parts that have an x with the other parts that have an x, and those with a y with the rest of those that also have a y. That's what "like terms" means :)

OpenStudy (turingtest):

It makes no difference the order of the expressions, it's just convention to put variables (letters) in the front, but that may need to change in order to solve a problem, so don't get locked into it.

OpenStudy (anonymous):

ok thank you